OpenStudy (anonymous):

when simplifying [3(x/3)+5]-15 i get x-10, is there any values of x that i should exclude from the composite function such that x cannot equal ?

OpenStudy (anonymous):

i'm sorry, i just get x

OpenStudy (sumeer):

no it's x-10. u musnt multiple 5 by 3. x-10 is ok.

OpenStudy (anonymous):

when i put it into wolframalpha it said x /: ahhhhh

OpenStudy (anonymous):

maybe u write it wrong in wolframalpha

OpenStudy (sumeer):

yes, i've copied and pasted and got x-10

OpenStudy (anonymous):

oh, ok. thank you. now, for the second part of the question, im given [(3x-15)/(3)]+5. i also got x for this one. would these be composite functions...if the first one is x-10 and the second one is x?

OpenStudy (anonymous):

and i did type it wrong into wolfrmalpha. it is supposed to be x for the first one. the 3 should be outside of the parentheses

OpenStudy (sumeer):

i think u mean f(x)=x-10 and g(x)=x then u ask (fog)(x), answer is x-10. i think u can do it.

OpenStudy (anonymous):

noo. the composites are f(x)=3x-15 and g(x)=(x/3)+5

OpenStudy (sumeer):

i just dont want to write answer, look at http://www.mathsisfun.com/sets/functions-composition.html then if u still cant come over it let me know

OpenStudy (anonymous):

so x for them isnt right when f0g and gof?

OpenStudy (precal):

there is not number that you need to exclude because f(x) and g(x) are linear functions

OpenStudy (anonymous):

and they both simplify to x, right?

OpenStudy (precal):

are they inverses of each other?

OpenStudy (precal):

not all compositions will simplify to x

OpenStudy (sumeer):

yeah x is right

OpenStudy (precal):